A hospice pharmacist receives a prescription for 1.1% morphine sulfate oral solution. She has a 60 mL bottle of morphine sulfate labeled 20 mg/5 mL and a large bottle of morphine labeled 100 mg/5 mL. How much of the 100 mg/5 mL product must be mixed with the content of the 20 mg/5 mL morphine sulfate bottle to prepare the desired percentage strength for the patient?
Answer #1

The percentage strength desired for the patient = 1.1 g morphine sulphate/100 mL solution

= 1100 mg morphine sulphate/100 mL solution

= 11 mg/mL

Now, {60 mL * (20 mg/5 mL) + x mL * (100 mg/5 mL)}/(60+x) mL = 11 mg/mL

i.e. (240 + 20x)/(60+x) = 11

i.e. 240 + 20x = 660 + 11x

i.e. 9x = 420

i.e. x = 46.7

Hence, 46.7 mL (~ 47 mL) of 100 mg/5 mL product must be mixed with the content (60 mL) of the 20 mg/5 mL morphine sulfate bottle to prepare the desired percentage strength for the patient (1.1%).